WHEN а) RiND Laplace transtorm (5 +e3t+56 2+) COS(6+) F(s) of bs find INVERSE Laplace transtorm ert) F(s) = +229 +5 2
Answer #1

a)

we have

L\left \{ (5+e^{3t}+5e^{-2t})cos(6t) \right \}=L\left \{ 5cos(6t)+e^{3t}cos(6t)+5e^{-2t}cos(6t) \right \}

L\left \{ (5+e^{3t}+5e^{-2t})cos(6t) \right \}=5L\left \{ cos(6t)\right \}+L\left \{ e^{3t}cos(6t) \right \}+5L\left \{e^{-2t}cos(6t) \right \}

we know that,

L\left \{ cos(at) \right \}=\frac{s}{s^2+a^2}

L\left \{ e^{bt}cos(at) \right \}=\frac{s-b}{(s-b)^2+a^2}

hence,

L\left \{ (5+e^{3t}+5e^{-2t})cos(6t) \right \}=5 \cdot \frac{s}{s^2+6^2}+\frac{s-3}{(s-3)^2+6^2}+5\cdot \frac{s-(-2)}{(s-(-2))^2+6^2}

L\left \{ (5+e^{3t}+5e^{-2t})cos(6t) \right \}=\frac{5s}{s^2+36}+\frac{s-3}{(s-3)^2+36}+\frac{5(s+2)}{(s+2)^2+36}

b)

we have

L^{-1}\left \{ F(s) \right \}=L^{-1}\left \{ \frac{2}{s^2+2s+5} \right \}

f(t)=L^{-1}\left \{ \frac{2}{s^2+2s+1+4} \right \}

f(t)=L^{-1}\left \{ \frac{2}{(s+1)^2+2^2} \right \}

we know that,

L^{-1}\left \{\frac{b}{(s+b)^2+a^2} \right \}=e^{-bt}sin(at)

hence

f(t)=e^{-t}sin(2t)